5. A projectile is fired vertically upwards with height h given by: h = 25t - 4.9^2 where 0 ≤ t ≤ 5


1
Expert's answer
2021-05-07T14:13:02-0400

if we need find:"h(5)=25*5-4.9*5*5=25(5-4.9)=2.5"

if we need find"h_{max},h'(t)=25-2*4.9*t" decreases,"h'(t_*)=0" ,

"t_*=\frac{25}{2*4.9}" ,"\implies h_{max}=h(t_*)=25*\frac{25}{2*4.9}-4.9*\frac{25}{2*4.9}*\frac{25}{2*4.9}="

"=\frac{625}{2*4.9}(1-\frac{1}{2})=\frac{625}{4*4.9}"

if we need find"t_{1,2}|h(t_{1,2})=0,h(t)=t*(25-4.9t)=0\implies t_1=0,t_2=\frac{25}{4.9}"